OpenStudy (anonymous):

a sheet of paper has a length of 29cm, a width of 8 in and a thickness of .10mm . What is the volume in cm^3

OpenStudy (anonymous):

@ganeshie8 pls help again sorry for the disturbance mr.

OpenStudy (unklerhaukus):

Volume is length times width times thickness\[V=\ell w t\] \[V=29[\text{cm}]\times8 [\text{in}]\times0.10[\text{mm}]\]

OpenStudy (unklerhaukus):

before the multiplication you should convert those units \[1[\text{in}]=2.45[\text{cm}]\] \[1[\text{mm}]=0.1[\text{cm}]\]

OpenStudy (unklerhaukus):

\[V=29[\text{cm}]\times8 \times2.45[\text{cm}]\times0.10\times0.1[\text{cm}]\\\quad =29\times8 \times2.45\times0.10\times0.1[\text{cm}][\text{cm}][\text{cm}]\\\quad=\qquad\dots\qquad [\text{cm}]^3\]

OpenStudy (anonymous):

what's the answer? sorry i really didnt understand

OpenStudy (anonymous):

@UnkleRhaukus

OpenStudy (unklerhaukus):

which bit is giving you trouble?

OpenStudy (anonymous):

the 0.1

OpenStudy (unklerhaukus):

there are 10 mm in 1 cm 1 mm in 0.1 cm

OpenStudy (unklerhaukus):

0.1mm = 0.01 cm

OpenStudy (unklerhaukus):

yeah?

OpenStudy (whpalmer4):

Actually, 1 in = 2.54 cm, not 2.45 cm... \[V=\ell w t\]\[V = 29 \text{ [cm]} * 8 \text{ [in]} * \frac{2.54 \text{ [cm]}}{1\text{ [in]}} * 0.10 \text{ [mm]} * \frac{1\text{ [cm]}}{10\text{ [mm]}}\]\[V = 29 \text{ [cm]} * 8 \cancel{\text{ [in]}} * \frac{2.54 \text{ [cm]}}{1\cancel{\text{ [in]}}} * 0.10 \cancel{\text{ [mm]}} * \frac{1\text{ [cm]}}{10\cancel{\text{ [mm]}}}\]\[V = 29 * 8 * 2.54 * 0.10 * \frac{1}{10} \text{ [cm]}^3 = \]

OpenStudy (unklerhaukus):

whoops.